OpenStudy (anonymous):

I am trying to do my homework but i lost my notes, how do i write an expression in a+bi ex1: [2(cos50˚+i sin50˚)]^3 ex2: (2-2i)^8 and once i have done that how do i find IIvII

OpenStudy (amistre64):

assuming you meant: a+bi this can be modeled as the point (a,b) to determine how to put it into polar form

OpenStudy (amistre64):

|v| is just the mod, or distance that (a,b) is from (0,0)

OpenStudy (amistre64):

mod(a+bi) = sqrt(a^2+b^2) arg(a+bi) = arctan(b/a)

OpenStudy (anonymous):

"this can be modeled as the point (a,b) to determine how to put it into polar form" how do i do this, I'm still confused :(

OpenStudy (amistre64):

if i give you a point (3,2) how would you put this onto a graph?

OpenStudy (amistre64):

the complex number: 3 + 2i can be modeled as the point (3,2)

OpenStudy (amistre64):



OpenStudy (amistre64):

the distance of the "point" from the "origin" gives you |v|

OpenStudy (amistre64):